OpenStudy (anonymous):

solve and round answer to 3 decimal places 53.2 = 2^x

OpenStudy (tkhunny):

Can you introduce a logarithm?

OpenStudy (anonymous):

\[x=\log_{2}53.2 \] thats what i was given

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

I see. You made it harder. Good work on your manipulation. You did that correctly, but that will not solve the problem. Try this: \(\log_{2}(53.2) = \dfrac{\log(52)}{\log(2)}\). That should open it up for a calculator with standard logarithm functions.